OpenStudy (mgriffin16):

Write the equation: One fifth of a number n is equal to -7. What is this as an equation?

OpenStudy (0487308):

\[n/5 = -7\]

OpenStudy (smartanne):

1/5n=-7

OpenStudy (0487308):

Both of these answers are correct, so you can use either one ;)

OpenStudy (mgriffin16):

Thank you

OpenStudy (0487308):

No problem.

OpenStudy (0487308):

Thanks for the medal :D

OpenStudy (welshfella):

though its best not to write the equation as 1/5n = -7 because it can be interpreted in 2 ways